Gabriella had to take three tests in her accounting class. She scored 71 points on her first test and 85 points on her second test. If her average for the three tests is 83, what score did Gabriella get on her third test?
Answer by stanbon(75887)

Gabriella had to take three tests in her accounting class. She scored 71 points on her first test and 85 points on her second test. If her average for the three tests is 83, what score did Gabriella get on her third test?
--------------
Average = (total points)/(# of tests)
----
83 = (71+85+x)/3
156+x = 3*83
------
x = 249-156
---
x = 93
-------
Cheers,
Stan H.